The length of the hypotenuse of an isosceles right-angled triangle is 12cm. What are the lengths of the other two sides?
Answer by mareebsiddiqui(7)

If it's an isoceles right triangle then it means that the The perpendicular and the base side are equal in length and let's suppose them to be 'x'.
So using the Pythagoras Theorem:
x^2 + x^2 = 12^2
4x^2 = 144
x^2 = 144/4 = 36
√x = √36 = 6